# Distribution of mixed character sums and extremal problems for Littlewood polynomials

### v1

@@ Abstract.
for fixed $\theta\in[0,1]$ and random character $\chi{\text{\rm\ (mod~$q$)}}$, as well as for a fixed character $\chi$ and randomly sampled $\theta\in[0,1].$ We present various applications of our results. For example, we construct Littlewood polynomials with large Mahler measure, thus establishing a new record in the Mahler problem (1963). We also show that $L_{2k}$ norms of well-known Turyn polynomials are asymptotically minimized at the shift $\alpha=1/4,$ proving a conjecture of Günther and Schmidt. An important ingredient in our work is a general way of dealing with “log-integrability” problems.
@@ 1. Introduction
Following Montgomery and Vaughan [MV-meanvalues], Bober and Goldmakher [BG-maximum] and later Bober, Goldmakher, Granville and Koukoulopoulos [BGGK] studied the distribution of $\max_{x}S(\chi,x,0)$ as $\chi$ varies over the characters mod $q$. Hussain [hussain-character-paths], inspired by [kowalski-sawin], considered the sum $S(\chi,tq,0)$ as a function of $t$ and studied its distribution for randomly chosen $\chi$. Subsequently the more difficult case of quadratic characters was dealt with by Lamzouri [lamzouri-quadraticmax] and Hussain and Lamzouri [lamzouri-hussain], and more recently Lamzouri and Nath [nath-lamzouri] study the distribution of the maximum of partial sums of cubic characters.
In the case when $\theta\neq 0$, Wang and Xu [wang-xu] show that a smoothed version of $S(\chi,x,\theta)$ is typically of size $\sqrt{x}$ for all irrational $\theta$ satisfying a weak Diophantine condition, contrasting Harper’s result. Very recently, Dell and Milićević [dell-milicevic] investigated the distribution of the incomplete Gauss sums $t\rightarrow S(\chi_{q},tq,1/q)$, where $\chi_{q}$ is the Jacobi symbol mod $q$ and $q$ is chosen randomly in a dyadic interval, which is an example of Problem 1 with rational $\theta\neq 0.$
In general, the limiting behavior significantly depends on the length $x=x(q)$ of the character sum, and in this paper we confine ourselves to the fixed range $\frac{x}{q}\to\beta\neq 0$ when $q\to\infty.$ To this end, we consider for a prime $q$ and a primitive (non-principal) Dirichlet character $\chi$ mod $q$ the shifted-mixed character sum
$S(\chi,\alpha,\beta,\theta)=\sum_{\alpha q<n<(\alpha+\beta)q}\chi(n)e(n\theta).$
@@
$\sum_{\alpha q<n<(\alpha+\beta)q}\chi(n)x^{n}$
@@
$F_{k,\chi,\alpha,\beta}(t)=\frac{e(-\alpha k)}{q^{1/2}}\sum_{\alpha q<n<(\alpha+\beta)q}\chi(n)e(n(k+t)/q),$
a normalised version of the sum $S(\chi,\alpha,\beta,\theta)$, with $\theta=(k+t)/q$ and $t\in[0,1).$ We regard $F_{k,q,\alpha,\beta}:\hat{\mathbb{F}}_{q}\rightarrow\mathscr{C}[0,1]$ as the discrete random process obtained by choosing $\chi$ uniformly at random from the set of all characters mod $q.$ We will consider the limiting behavior of $F_{k,q,\alpha,\beta}$ as $q\rightarrow\infty$ over the primes, with $k$ either fixed or tending to infinity as well.
@@
where $\mathbb{W}(l)$ is a completely multiplicative Steinhaus random variable and $\mathbb{X}$ is a random variable uniformly distributed on the unit circle, independent of $\mathbb{W}$. As we shall see, when $k$ tends to infinity the effect of multiplicativity vanishes, and so we also define the random process
@@
where the $\mathbb{X}(l)$ are independent realizations of $\mathbb{X}$. Both of these processes are almost surely continuous. Our first main result is the following convergence statement in the space of continuous functions $\mathscr{C}[0,1].$
@@
Fix real numbers $\alpha$ and $\beta$. Then
@@
$\frac{1}{q^{1/2}}\sum_{n<\beta q}\chi(n)e(n\theta)\xrightarrow{\phantom{dd}d\phantom{dd}}F_{0,\beta}(t_{0})$
@@
We now turn to Problem 2 and fix $\chi_{q}$ as the quadratic character mod $q$. We partition the interval $[0,1]$ into pieces $\frac{k}{q},\frac{k+1}{q})$ for $k\in[0,q-1]$ and our aim is to show that the behavior of the character sum $S(\chi,\alpha,\beta,\theta)$ in a randomly chosen subinterval is governed by an explicit random point process. For each prime $q$, let $G_{q,\alpha,\beta}:\{0,1,\ldots,q-1\}\rightarrow\mathscr{C}[0,1]$ be the random process $k\rightarrow\frac{q^{1/2}}{\tau(\chi_{q})}F_{k,\chi_{q},\alpha,\beta}(t)$, where the set $\{0,1,\ldots,q-1\}$ is equipped with the uniform measure. Finally, we introduce the random process
@@ 1.1. Application to the Mahler problem
Our first application concerns a classical quantity of an algebraic polynomial $P\in\mathbb{C}[x].$ Recall the $L_{\lambda}$ norm of $P$, given, for $\lambda>0$, by
@@
(1) $M_{0}(P):=\exp\left(\int_{0}^{1}\log|P(e(t))|\mathrm{d}t\right).$
@@
Problem 3 is a younger cousin of the infamous Lehmer problem [Lehmer], asking for integer polynomials with a small Mahler measure. It easy to see that $\frac{M_{0}(P)}{\|P\|_{2}}<1,$ but even the question of whether $b_{m}=1$ is wide open. The conclusion that $b_{m}=1$ would immediately follow from the existence of “ultra-flat” polynomials (Littlewood conjecture) with coefficients $\pm 1.$ Indeed, if $P_{n}$ is a Littlewood polynomial satisfying the pointwise bound
@@
Interestingly, in $1970$, Fielding [Field] showed directly (without appealing to the ultra-flatness) that
@@
The Mahler problem, which may be regarded as a “flatness” question — albeit in a somewhat weaker form — has attracted considerable attention over the last few decades. Choi and Erdélyi [CE] proved that for each $n$, there exists Littlewood polynomial $f_{n}$ satisfying $M_{0}(f_{n})/\sqrt{n}\geq 1/2+o(1)$ and consequently $b_{m}\geq 1/2.$ In [CE1] they established a stronger result, determining the expected value of the normalized Mahler measure for random $\pm 1$ Littlewood polynomials
@@
There exist Littlewood polynomials of arbitrarily large degree with normalized Mahler measure >.954>.954. Consequently, $b_{m}>.954$.
@@
$\exp\left(\int_{0}^{1}\log\left|\frac{1}{(1.1q)^{1/2}}S(\chi_{q},.2,1.1,\theta)\right|\mathrm{d}\theta\right)\longrightarrow c\approx.954\ldots$
@@
These values come from an approximate numerical calculation, and we do not wish to give the impression that the Mahler measure is maximized at these precise values of $\alpha$ and $\beta$. The flexibility of our methods, however, offers a possible way to modify the construction of the polynomials further, obtaining better bounds. We will explore this in future work.
In fact, Mossinghoff [mossinghoff-mahler] already conjectured that a larger normalized Mahler measure might be achievable by considering $\beta\neq 1$, in analogy with the result of Jedwab, Katz, and Schmidt [jedwab] that the merit factor of the generalized Turyn polynomial is maximized at $\alpha\approx.221$ and $\beta=1.058$. Using methods of this paper one can show that for these values
$\exp\left(\int_{0}^{1}\log\left|\frac{1}{(1.058q)^{1/2}}S(\chi_{q},.221,1.058,\theta)\right|\mathrm{d}\theta\right)\approx.9535,$
which also provides improvement over the bound in [mossinghoff-mahler] and shows that for the generalized Turyn polynomials, the merit factor and the normalized Mahler measure are not maximized at the same point.
1.2. Application to the Günther-Schmidt conjecture
MIn 2017, Günther and Schmidt [GuentherSchmidt2017] determined the limiting value of the normalized $L_{2k}$ norm of the Fekete and Turyn polynomials when $k\in\mathbb{N}$. More precisely, they proved that there is a function $\phi_{k}:\mathbb{R}\to\mathbb{R}$ such that $\lim_{q\to\infty}\|\frac{F_{q,\left\lfloor\alpha q\right\rfloor}}{\sqrt{q}}\|_{2k}=\phi_{k}(\alpha)$ and described a method to compute $\phi_{k}(0).$ The expressions for $\phi_{k}(\alpha)$ are given by a rather complicated recursive combinatorial identities, which make them hard to use in practice. However, they showed that $\phi_{k}(\alpha)$ attains its minimum at $\alpha=\frac{1}{4}$ for $k=2,3,4$ and conjectured that this continues to hold for every $k\geq 5.$ As an application of our results, we confirm this conjecture.
@@ 1.2. Application to the Günther-Schmidt conjecture
$\emph{argmin}_{\alpha\in[0,1]}\phi_{k}(\alpha)=\frac{1}{4}.$
Theorem 4 implies that the polynomials $F_{q,\left\lfloor\frac{q}{4}\right\rfloor}(\theta):=S(\chi_{q},\frac{1}{4},1,\theta)$ tend to be somewhat flat, and therefore it is natural to expect that the Mahler measure $M_{0}(F_{q,\left\lfloor\frac{q}{4}\right\rfloor})$ is large, explaining the large value achieved in [mossinghoff-mahler].
@@ 2. Roadmap to the proofs
From this formula we begin to see how the random processes that appear in Theorems 1 and 2 emerge. Indeed, for most characters, the sum (2) is well approximated by a short sum and so we expect to establish convergence to the random process
@@
As we shall see, for large $k=k_{q}\to\infty$, the multiplicativity generally becomes unimportant at the cost of removing a zero density subset of primes, so that the limiting process becomes
@@
where $k=k_{q}$ and the variables $m_{i}$ and $n_{j}$ are small (depending on $q$). One would like to argue that the mod $q$ condition becomes irrelevant as $q$ gets large, so that the moments of (2) coincide with the random multiplicative model. This is indeed what happens for fixed $k$. The mod $q$ condition also becomes irrelevant when $k_{q}=q^{o(1)}$, and in this case the moments of (2) coincide with the random model without multiplicativity. However, for certain values of $k=k_{q}\to\infty$, this need not be the case; if, say, $k\equiv(q+1)/3\pmod{q}$ (or $k=\left\lfloor q\theta\right\rfloor$ is very close to $(q+1)/3$ for irrational $\theta$), then this equation becomes
@@
and the mod $q$ condition does indeed become superfluous by size considerations, but this shows that the mod $q$ condition was an essential feature of the first equation. In the case $k_{q}=\left\lfloor q\theta\right\rfloor$, such linear coincidences happen only finitely often for any irrational $\theta$, however, so do not present a problem.
More insidious are the coincidences of a higher degree. .For any fixed choice of $m_{1},\ldots,m_{d}$ and $n_{1},\ldots,n_{d}$, equation (3) becomes a degree d1d-1 equation in $k$, so we naturally arrive at the question of how often is $k=k_{q}$ the solution of polynomial equations with small coefficients mod $q.$ In Lemma 6, we show that when $k_{q}=o(\pi(q))$ or $k_{q}=\left\lfloor q\theta\right\rfloor$ for irrational $\theta$, this happens for $o(\pi(X))$ choices of primes $q\leq X$. We then use iterative density arguments to construct a density one set of suitable primes $\mathcal{P}_{\theta}$ to avoid such coincidences for higher moments when $q\to\infty.$
We now describe the key ideas to prove Theorem 3. We write the logarithmic Mahler measure of $S_{q}:=S(\chi_{q},\alpha,\beta,\theta)$ as
$\log M_{0}(S_{q})=\int_{0}^{1}\log\left|q^{1/2}\frac{S_{q}(e(t))}{q^{1/2}}\right|dt=\log\sqrt{q}+\frac{1}{q}\sum_{k=0}^{q-1}\int_{0}^{1}\log\bigg|F_{q,\alpha,\beta}\bigg(e\bigg(\frac{k+t}{q}\bigg)\bigg)\bigg|\mathrm{d}t.$
@@
$\frac{1}{q}\sum_{k=0}^{q-1}\int_{0}^{1}\log\bigg|F_{q,\alpha,\beta}\bigg(e\bigg(\frac{k+t}{q}\bigg)\bigg)\bigg|\mathrm{d}t\longrightarrow\operatorname{\mathbb{E}}\left[\int_{0}^{1}\log\left|G_{\alpha,\beta}(t)\right|\mathrm{d}t\right].$
However, the functional $\ell(f)=\int_{0}^{1}\log|f(t)|dt$ is not continuous on $\mathscr{C}[0,1]$ and so we cannot immediately apply Theorem 2. We thus have to deal with the issues of uniform log-integrability, which is a difficult problem for both random and deterministic series. To this end, we consider the regularized continuous functional
$\ell_{\varepsilon}(f)=\int_{0}^{1}\log(|f(t)|)w_{\varepsilon}(\left|f(t)\right|)\mathrm{d}t$
with $w_{\varepsilon}(t)$ a smooth minorant of the indicator function $\mathds{1}_{t\geq\varepsilon}$, to which our convergence results of Theorem 2 apply.
@@
Finally, in order to prove Theorem 4, we use our Theorem 2 to conveniently rewrite the norms $L_{2k}$ and then establish a relevant monotonicity property of the modified Lerch transcendent to arrive at the result.

### v2

@@ Abstract.
for fixed $\theta\in[0,1]$ and random character $\chi{\text{\rm\ (mod~$q$)}}$, as well as for a fixed character $\chi$ and randomly sampled $\theta\in[0,1].$ We present various applications of our results. For example, we construct Littlewood polynomials with large Mahler measure and $L_{1}$ norm, thus establishing new records in the Mahler and Newman problems. We also show that $L_{2k}$ norms of well-known Turyn polynomials are asymptotically minimized at the shift $\alpha=1/4,$ proving a conjecture of Günther and Schmidt. An important ingredient in our work is a general way of dealing with “log-integrability” problems.
@@ 1. Introduction
Following Montgomery and Vaughan [MV-meanvalues], Bober and Goldmakher [BG-maximum] and later Bober, Goldmakher, Granville and Koukoulopoulos [BGGK] studied the distribution of $\max_{x}S(\chi,x,0)$ as $\chi$ varies over the characters mod $q$. Hussain [hussain-character-paths], inspired by [kowalski-sawin], considered the sum $S(\chi,tq,0)$ as a function of $t$ and studied its distribution for randomly chosen $\chi$. Subsequently the more difficult case of quadratic characters was dealt with by Lamzouri [lamzouri-quadraticmax] and Hussain and Lamzouri [lamzouri-hussain], and more recently Lamzouri and Nath [nath-lamzouri] studied the distribution of the maximum of partial sums of cubic characters.
In the case when $\theta\neq 0$, Wang and Xu [wang-xu] showed that a smoothed version of $S(\chi,x,\theta)$ is typically of size $\sqrt{x}$ for all irrational $\theta$ satisfying a weak Diophantine condition, contrasting Harper’s result. Very recently, Dell and Milićević [dell-milicevic] investigated the distribution of the incomplete Gauss sums $t\rightarrow S(\chi_{q},tq,1/q)$, where $\chi_{q}$ is the Jacobi symbol mod $q$ and $q$ is chosen randomly in a dyadic interval, which is an example of Problem 1 with rational $\theta\neq 0.$
In general, the limiting behavior significantly depends on the length $x=x(q)$ of the character sum, and in this paper we confine ourselves to the fixed range $\frac{x}{q}\to\beta\neq 0$ when $q\to\infty.$ To this end, we consider for a prime $q$ and a primitive (non-principal) Dirichlet character $\chi$ mod $q$ the shifted mixed character sum
$S(\chi,\alpha,\beta,\theta)=\sum_{\alpha q<n\leq(\alpha+\beta)q}\chi(n)e(n\theta).$
@@
$\sum_{\alpha q<n\leq(\alpha+\beta)q}\chi(n)x^{n}$
@@
$F_{k,\chi,\alpha,\beta}(t)=\frac{e(-\alpha k)}{q^{1/2}}\sum_{\alpha q<n\leq(\alpha+\beta)q}\chi(n)e(n(k+t)/q),$
a normalised version of the sum $S(\chi,\alpha,\beta,\theta)$, with $\theta=(k+t)/q$ and $t\in[0,1).$ We regard $F_{k,q,\alpha,\beta}:\widehat{\mathbb{F}}_{q}\rightarrow\mathscr{C}[0,1]$ as the (discrete) random process obtained by choosing $\chi$ uniformly at random from the set of all characters mod $q.$ We will consider the limiting behavior of $F_{k,q,\alpha,\beta}$ as $q\rightarrow\infty$ over the primes, with $k$ either fixed or tending to infinity as well.
@@
where $\mathbb{W}(l)$ is a completely multiplicative Steinhaus random variable and $\mathbb{X}$ is a random variable uniformly distributed on the unit circle, independent of $\mathbb{W}$. As we shall see, when $k$ tends to infinity the effect of multiplicativity subsides, and so we also define the random process
@@
where the $\mathbb{X}(l)$ are independent realizations of $\mathbb{X}$. Both of these processes are almost-surely continuous. Our first main result is the following convergence statement in the space of continuous functions $\mathscr{C}[0,1].$
@@
Fix real numbers $\alpha$ and $\beta$. Then:
@@
$\frac{1}{q^{1/2}}\sum_{n\leq\beta q}\chi(n)e(n\theta)\xrightarrow{\phantom{dd}d\phantom{dd}}F_{0,\beta}(t_{0})$
@@
We now turn to Problem 2 and fix $\chi_{q}$ as the quadratic character mod $q$. We partition the interval $[0,1]$ into pieces $\frac{k}{q},\frac{k+1}{q})$ for $k\in[0,q-1]$ and our aim is to show that the behavior of the character sum $S(\chi,\alpha,\beta,\theta)$ in a randomly chosen subinterval is governed by an explicit random process.
For each prime $q$, let $G_{q,\alpha,\beta}:\{0,1,\ldots,q-1\}\rightarrow\mathscr{C}[0,1]$ be the random process $k\rightarrow\frac{q^{1/2}}{\tau(\chi_{q})}F_{k,\chi_{q},\alpha,\beta}(t)$, where the set $\{0,1,\ldots,q-1\}$ is equipped with the uniform measure. Finally, we introduce the random process
@@ 1.1. Application to the Mahler problem
Our first application concerns a classical quantity of a polynomial $P\in\mathbb{C}[x].$ Recall the $L_{\lambda}$ norm of $P$, given, for $\lambda>0$, by
@@
(1) $M_{0}(P)=\exp\left(\int_{0}^{1}\log|P(e(t))|\mathrm{d}t\right).$
@@
Problem 3 is a younger cousin of the infamous Lehmer problem [Lehmer], asking for integer polynomials with a small Mahler measure. It easy to see that $\frac{M_{0}(P)}{\|P\|_{2}}<1,$ but even the question of whether $b_{m}=1$ is wide open. The conclusion that $b_{m}=1$ would immediately follow from the existence of “ultra-flat” polynomials (Littlewood’s conjecture) with coefficients $\pm 1.$ Indeed, if for every $\varepsilon>0$ there is a Littlewood polynomial $P_{n}$ satisfying the pointwise bound
@@
Interestingly, in $1970$, Fielding [Field] showed directly (without appealing to ultra-flatness) that
@@
The Mahler problem, which may be regarded as a “flatness” question — albeit in a somewhat weaker form — has attracted considerable attention over the last few decades. Choi and Erdélyi [CE] proved that for each $n$, there exists a Littlewood polynomial $f_{n}$ satisfying $M_{0}(f_{n})/\sqrt{n}\geq 1/2+o(1)$ and consequently $b_{m}\geq 1/2.$ In [CE1] they established a stronger result, determining the expected value of the normalized Mahler measure for random $\pm 1$ Littlewood polynomials, namely
@@
There exist Littlewood polynomials of arbitrarily large degree with normalized Mahler measure >0.954>0.954. Consequently, $b_{m}>0.954$.
@@
$\frac{1}{(1.1q)^{1/2}}\exp\left(\int_{0}^{1}\log\left|S(\chi_{q},0.2,1.1,\theta)\right|\mathrm{d}\theta\right)\longrightarrow c\approx 0.954\ldots$
@@
The values of $\alpha$ and $\beta$ come from an approximate numerical calculation, and we do not wish to give the impression that the Mahler measure is maximized at these precise values. The flexibility of our methods, however, offers a possible way to modify the construction of the polynomials further, obtaining better bounds. We will explore this in future work.
In fact, Mossinghoff [mossinghoff-mahler] already conjectured that a larger normalized Mahler measure might be achievable by considering $\beta\neq 1$, in analogy with the result of Jedwab, Katz and Schmidt [jedwab] that the merit factor of the generalized Turyn polynomial is maximized at $\alpha\approx 0.221$ and $\beta=1.058$. Using methods of this paper one can show that for these values
$\frac{1}{(1.058q)^{1/2}}\exp\left(\int_{0}^{1}\log\left|S(\chi_{q},0.221,1.058,\theta)\right|\mathrm{d}\theta\right)\approx 0.9535,$
which also provides an improvement over the bound in [mossinghoff-mahler] and shows that for the generalized Turyn polynomials, the merit factor and the normalized Mahler measure are not maximized at the same point.
Similar extremal problems have been considered for other norms, most notably for the $L_{1}$ norm, which is known as Newman’s problem. Mossinghoff records the new lower bound of $0.9775\ldots$ for the ratio $P_{n}||_{1}/\sqrt{n}$ for an infinite family of Littlewood polynomials. Following our methods, this is improved slightly to $0.9783\ldots$ by the generalized Turyn polynomials with parameters $\alpha=0.2$ and $\beta=1.1$.
1.2. Application to the Günther–Schmidt conjecture
MIn 2017, Günther and Schmidt [GuentherSchmidt2017] determined the limiting value of the normalized $L_{2k}$ norm of the Fekete and Turyn polynomials when $k\in\mathbb{N}$. More precisely, they proved that there is a function $\phi_{k}:\mathbb{R}\to\mathbb{R}$ such that $\lim_{q\to\infty}\frac{1}{{\sqrt{q}}}\|S(\chi_{q},\alpha,1,\cdot)\|_{2k}=\phi_{k}(\alpha)$ and described a method to compute $\phi_{k}(0).$ The expressions for $\phi_{k}(\alpha)$ are given by rather complicated recursive combinatorial identities, which make them hard to use in practice. However, they showed that $\phi_{k}(\alpha)$ attains its minimum at $\alpha=\frac{1}{4}$ for $k=2,3,4$ and conjectured that this continues to hold for every $k\geq 5.$ As an application of our results, we confirm this conjecture.
@@ 1.2. Application to the Günther–Schmidt conjecture
$\operatorname{argmin}_{\alpha\in[0,1]}\phi_{k}(\alpha)=\frac{1}{4}.$
Theorem 4 implies that the polynomials $S(\chi_{q},\frac{1}{4},1,\theta)$ tend to be somewhat flat, and therefore it is natural to expect that the Mahler measure $M_{0}(S(\chi_{q},\frac{1}{4},1,\cdot))$ is large, explaining the large value achieved in [mossinghoff-mahler].
@@ 2. Roadmap to the proofs
From this formula we begin to see how the random processes that appear in Theorems 1 and 2 emerge. Indeed, for most characters, the sum (2) is well approximated by a short sum, and so via the method of moments we will establish convergence to the random process
@@
As we shall see, for large $k=k_{q}\to\infty$, multiplicativity generally becomes unimportant at the cost of removing a zero density subset of primes, so that the limiting process becomes
@@
where $k=k_{q}$ and the variables $m_{i}$ and $n_{j}$ are small (relative to $q$). One would like to argue that the mod $q$ condition becomes irrelevant as $q$ gets large, so that the moments of (2) coincide with the random multiplicative model $F_{k,\alpha,\beta}$. This is indeed what happens for fixed $k$. The mod $q$ condition also becomes irrelevant when $k_{q}=q^{o(1)}$, and in this case the moments of (2) coincide with the random model $F_{\alpha,\beta}$ without multiplicativity. However, for certain values of $k=k_{q}\to\infty$, this need not be the case; if $k\equiv(q+1)/3\pmod{q}$ (or $k=\left\lfloor q\theta\right\rfloor$ is very close to $(q+1)/3$ for irrational $\theta$), say, then this equation becomes
@@
and the mod $q$ condition does indeed become superfluous by size considerations, but this shows that the mod $q$ condition was an essential feature of the first equation (3). In the case $k_{q}=\left\lfloor q\theta\right\rfloor$, such linear coincidences happen only finitely often for any irrational $\theta$, however, so do not present a problem.
More insidious are the coincidences of a higher degree. .For any fixed choice of $m_{1},\ldots,m_{d}$ and $n_{1},\ldots,n_{d}$, the equation (3) becomes a degree d1d-1 equation in $k$, so we naturally arrive at the question of how often is $k=k_{q}$ the solution of polynomial equations with small coefficients mod $q.$ In Lemma 6, we show that when $k_{q}=o(\pi(q))$ or $k_{q}=\left\lfloor q\theta\right\rfloor$ for irrational $\theta$, this happens for $o(\pi(X))$ choices of primes $q\leq X$. We then use iterative density arguments to construct a density one set of suitable primes $\mathcal{P}_{\theta}$ to avoid such coincidences for higher moments when $q\to\infty.$
We now describe the key ideas to prove Theorem 3. We write the logarithmic Mahler measure of $S_{q}(\theta):=S(\chi_{q},\alpha,\beta,\theta)$ as
$\log M_{0}(S_{q})=\int_{0}^{1}\log\left|q^{1/2}\frac{S_{q}(\theta)}{q^{1/2}}\right|\mathrm{d}\theta=\log\sqrt{q}+\frac{1}{q}\sum_{k=0}^{q-1}\int_{0}^{1}\log\bigg|F_{k,\chi_{q},\alpha,\beta}(t)\bigg|\mathrm{d}t.$
@@
$\frac{1}{q}\sum_{k=0}^{q-1}\int_{0}^{1}\log\bigg|F_{k,\chi_{q},\alpha,\beta}(t)\bigg|\mathrm{d}t\longrightarrow\operatorname{\mathbb{E}}\left[\int_{0}^{1}\log\left|G_{\alpha,\beta}(t)\right|\mathrm{d}t\right].$
Note, however, that the functional $\ell(f)=\int_{0}^{1}\log|f(t)|dt$ is not continuous on $\mathscr{C}[0,1]$ and so we cannot immediately apply Theorem 2. We thus have to deal with the issues of uniform log-integrability, which is a difficult problem for both random and deterministic series. To this end, we consider the regularized continuous functional
$\ell_{\varepsilon}(f)=\int_{0}^{1}\log\left|f(t)\right|w_{\varepsilon}(\left|f(t)\right|)\mathrm{d}t$
with $w_{\varepsilon}(t)$ a smooth minorant of the indicator function $\mathds{1}_{t\geq\varepsilon}$, to which our convergence results of Theorem 2 apply more easily.
@@
Finally, in order to prove Theorem 4, we use our Theorem 2 to conveniently rewrite the $L_{2k}$ norms and then establish a relevant monotonicity property of the modified Lerch transcendent to arrive at the result.